# Efficient Speech Language Modeling via Energy Distance in Continuous Latent Space

### v1

@@ 1 Introduction
Unlike text sequences composed of discrete tokens from a finite vocabulary, speech audio is often represented as a lengthy sequence of sampling points usually stored as integers or floating-point values within a range. This fundamental difference poses significant challenges for autoregressive speech modeling in a manner analogous to text. Text language models typically rely on a finite vocabulary to acquire a categorical distribution at each step, which is essential for stable training via cross-entropy loss and efficient ancestral sampling during inference. To adapt autoregressive modeling to speech, researchers have explored discretizing speech into sequences of discrete tokens using external quantization modules. Examples include discretization on SSL-learned features [24] in [32, 83, 48], discretization on ASR-supervised features in [7, 8, 81, 30] and reconstruction-oriented codec [80, 9, 31, 84, 78, 50, 25] in [1, 71, 86, 10, 79, 37]. However, discretization inevitably creates an information bottleneck, potentially losing the rich details in the raw waveform and thus degrading the reconstruction quality derived from the discretized tokens. One popular approach to mitigating this information loss is residual vector quantization [RVQ; 69, 80], which converts raw waveforms into multi-stream sequences of discrete tokens. However, RVQ introduces additional complexity, requiring hierarchical autoregressive architectures to effectively model multi-stream sequences [71, 76].
@@
In this research, we introduce an approach to speech language modeling in continuous latent space using energy distance (SLED). We model the per-step continuous distribution with a lightweight implicit conditional generative module that takes features encoded with autoregressive dependencies and random noise as inputs [36, 59]. To measure the discrepancy between the underlying data distribution and the model’s distribution, we employ a specialized form of maximum mean discrepancy [MMD; 17], known as generalized energy distance [GED; 41]. By selecting an appropriate distance function, the resulting metric becomes a strictly proper scoring rule [15, 58], enabling efficient training of the model to capture the autoregressive continuous distribution. Our approach eliminates the need for iterative sampling at each autoregressive step [36, 68] or post-AR refinement [45], significantly improving efficiency in modeling lengthy speech sequences while maintaining sufficient capacity. We demonstrate the effectiveness of our method in zero-shot and streaming speech synthesis, showing its strong potential for application in general-purpose speech language models.
@@ 3.3 Classifier-free Guidance
Continuous latent generation exhibits higher noise levels than its discrete counterpart. To improve the generation quality and the alignment with prompts, we adopt the classifier-free guidance technique [CFG; 22, 59] during inference. At each step of autoregressive generation, we perform an additional forward pass through $\psi$, during which the prompt text is masked out, to obtain $\bm{z}^{\prime}_{t}$. We then linearly interpolate $\bm{z}_{t}$ with $\bm{z}^{\prime}_{t}$ and use the result as input to the per-token generative module $g$:
@@ 4 Related Work
Given great success of autoregressive language modeling approach in text generation, researchers have started exploring whether speech sequences can be modeled in a similar way. These studies begin with speech synthesis as a foundational task [32, 1, 71], eventually extending to the development of general-purpose speech language models [47, 83, 10, 81, 37, 30]. [32] was the first to propose discretizing self-supervised representations of speech as semantic tokens and to perform autoregressive speech generation. The method of learning semantic tokens has since evolved to quantize ASR-supervised representations [7, 8]. Due to information loss, waveforms reconstructed from coarse semantic tokens are often unsatisfactory, necessitating modeling with acoustic tokens produced by reconstruction-oriented audio codecs [80, 9, 31, 84, 78, 50, 25]. Since the acoustic tokens generated by codecs often consist of multiple sequences from residual quantization, a hierarchical architecture is required for autoregressive speech modeling. [71] utilize an autoregressive model to generate the sequence of tokens from the first codebook, followed by a non-autoregressive model to generate tokens from the residual codebooks. To mitigate the latency introduced by post-AR refinement, [76, 86] apply the RQ-Transformer [34], which uses a smaller nested Transformer to model tokens from multiple codebooks within a single autoregressive step. Nonetheless, modeling multi-stream acoustic tokens introduces additional complexity. Inspired by advancements in image generation [67, 36], researchers have begun exploring performing speech language modeling in the continuous latent space. [45] use mel-spectrograms as the latent representation and model the per-step distribution with a regression loss. However, using a regression loss fails to capture the underlying distribution, requiring a post-AR refinement network and handcrafted repetition loss. [68] adopt a per-token diffusion loss [36], which requires iterative sampling for each autoregressive step, leading to significant inference latency. [72] models each step via flow matching [39], which also encounters similar problems. [85, 38] models the data at each autoregressive step with a Gaussian distribution or GMM and employs a VAE to obtain the target distribution parameters for each step. This approach imposes overly strong constraints on the latent space, limiting the model’s expressive capacity.
@@ 5.3 Evaluation Settings
We evaluate zero-shot speech synthesis performance using LibriSpeech test-clean set, ensuring that none of the test speakers are included in the training data. Following [71, 45], we use samples with durations between 4 and 10 seconds, resulting in a 2.2-hour subset comprising 1,234 samples and 40 unique speakers. Since the LibriSpeech consists of case-insensitive text without punctuation, while the expected input for models trained on LibriHeavy is unnormalized, we use the test-clean set of LibriSpeech-PC [44] to evaluate SLED. The LibriSpeech-PC is nearly identical to LibriSpeech, but its text has been restored to include capitalization and punctuation. [^3] The test-clean set of Librispeech-PC excludes a small number of low-quality samples (203 out of 2,620) from the original Librispeech test-clean set, thus the test subset after applying the same protocol contains only 1,154 samples. The evaluation is conducted under two settings: 1) 3s Prefix as Prompt: The model is provided with the text transcription and the first 3 seconds of the utterance as the prompt, and it is expected to perform speech continuation. 2) Reference Utterance as Prompt: The model is given a reference utterance from the same speaker along with its transcription as the prompt. Using the text of the target utterance, the model is expected to synthesize speech that retains the characteristics of the reference speaker. For this experiment, we follow the configuration outlined in [5]. We first filter the samples in LibriSpeech test-clean set based on length and order them by sample ID. For each speaker, the $i$-th speech sample is synthesized using the $(i-1)$-th sample as the prompt, while the first sample is synthesized using the last sample from the same speaker as the prompt. For Offline and Streaming Inference experiments, we follow previous data settings and no prompt speech is provided. We primarily use the following metrics to assess the intelligibility and in-context learning capability.
@@ 5.4 Main Results
Zero-shot TTS  Table 1 presents the results of zero-shot TTS. In both the speech continuation and reference utterance prompting tasks, SLED achieves a word accuracy surpassing the ground truth (1.59/1.51 vs. 1.78 in WER-C), demonstrating its modeling ability. Interestingly, we found that using a reference utterance as a prompt results in more accurate synthesis compared to using the prefix, showing the ability of in-context learning. Comparisons of SLED performance across different data scales are provided in App. A. We found that 1000 hours of speech is sufficient to achieve most of the generation and in-context learning capabilities. Further scaling the training data to 50,000 hours enhances these abilities even more. Notably, a comparison between SLED and VALL-E highlights key differences in autoregressive modeling approaches. Both models use Encodec as the latent encoder, but SLED performs modeling in the continuous domain, whereas VALL-E operates in the discrete domain. VALL-E-like hierarchical models require an extra non-autoregressive local model ($\approx$159M) to predict residual tokens. In contrast, SLED relies on a lightweight generative module ($\approx$35M) for continuous vector generation, achieving better parameter efficiency. Despite this, SLED outperforms VALL-E on all metrics. These results show the superiority of our continuous speech language modeling approach. Compared with other continuous approaches, SLED achieves better or similar quality while enjoys the advantage in generation efficiency. MELLE incorporates an NAR refinement, whereas SLED is purely autoregressive and can generate in a streaming manner. FELLE must perform multi-step integration iterations along a predefined ODE path at each decoding step, while SLED only requires a single forward pass through its generative module. Moreover, at roughly the same parameter scale, current speech language modeling approaches still exhibit a significant gap in voice cloning compared with traditional TTS models. However, Llasa [79] shows that scaling up can dramatically strengthen discrete autoregressive speech models, which inspires us to further explore the scaling capabilities of continuous autoregressive modeling approaches.

### v2

@@ 1 Introduction
Unlike text sequences composed of discrete tokens from a finite vocabulary, speech audio is often represented as a lengthy sequence of sampling points usually stored as integers or floating-point values within a range. This fundamental difference poses significant challenges for autoregressive speech modeling in a manner analogous to text. Text language models typically rely on a finite vocabulary to acquire a categorical distribution at each step, which is essential for stable training via cross-entropy loss and efficient ancestral sampling during inference. To adapt autoregressive modeling to speech, researchers have explored discretizing speech into sequences of discrete tokens using external quantization modules. Examples include discretization on SSL-learned features [24] in [33, 85, 49], discretization on ASR-supervised features in [7, 8, 83, 31] and reconstruction-oriented codec [82, 9, 32, 86, 80, 51, 25] in [1, 72, 88, 10, 81, 38]. However, discretization inevitably creates an information bottleneck, potentially losing the rich details in the raw waveform and thus degrading the reconstruction quality derived from the discretized tokens. One popular approach to mitigating this information loss is residual vector quantization [RVQ; 70, 82], which converts raw waveforms into multi-stream sequences of discrete tokens. However, RVQ introduces additional complexity, requiring hierarchical autoregressive architectures to effectively model multi-stream sequences [72, 77].
@@
In this research, we introduce an approach to speech language modeling in continuous latent space using energy distance (SLED). We model the per-step continuous distribution with a lightweight implicit conditional generative module that takes features encoded with autoregressive dependencies and random noise as inputs [37, 60]. To measure the discrepancy between the underlying data distribution and the model’s distribution, we employ a specialized form of maximum mean discrepancy [MMD; 17], known as generalized energy distance [GED; 42]. By selecting an appropriate distance function, the resulting metric becomes a strictly proper scoring rule [15, 59], enabling efficient training of the model to capture the autoregressive continuous distribution. Our approach eliminates the need for iterative sampling at each autoregressive step [37, 69] or post-AR refinement [46], significantly improving efficiency in modeling lengthy speech sequences while maintaining sufficient capacity. We demonstrate the effectiveness of our method in zero-shot and streaming speech synthesis, showing its strong potential for application in general-purpose speech language models.
@@ 3.3 Classifier-free Guidance
Continuous latent generation exhibits higher noise levels than its discrete counterpart. To improve the generation quality and the alignment with prompts, we adopt the classifier-free guidance technique [CFG; 22, 60] during inference. At each step of autoregressive generation, we perform an additional forward pass through $\psi$, during which the prompt text is masked out, to obtain $\bm{z}^{\prime}_{t}$. We then linearly interpolate $\bm{z}_{t}$ with $\bm{z}^{\prime}_{t}$ and use the result as input to the per-token generative module $g$:
@@ 4 Related Work
Given great success of autoregressive language modeling approach in text generation, researchers have started exploring whether speech sequences can be modeled in a similar way. These studies begin with speech synthesis as a foundational task [33, 1, 72], eventually extending to the development of general-purpose speech language models [48, 85, 10, 83, 38, 31]. [33] was the first to propose discretizing self-supervised representations of speech as semantic tokens and to perform autoregressive speech generation. The method of learning semantic tokens has since evolved to quantize ASR-supervised representations [7, 8]. Due to information loss, waveforms reconstructed from coarse semantic tokens are often unsatisfactory, necessitating modeling with acoustic tokens produced by reconstruction-oriented audio codecs [82, 9, 32, 86, 80, 51, 25]. Since the acoustic tokens generated by codecs often consist of multiple sequences from residual quantization, a hierarchical architecture is required for autoregressive speech modeling. [72] utilize an autoregressive model to generate the sequence of tokens from the first codebook, followed by a non-autoregressive model to generate tokens from the residual codebooks. To mitigate the latency introduced by post-AR refinement, [77, 88] apply the RQ-Transformer [35], which uses a smaller nested Transformer to model tokens from multiple codebooks within a single autoregressive step. Nonetheless, modeling multi-stream acoustic tokens introduces additional complexity. Inspired by advancements in image generation [68, 37], researchers have begun exploring performing speech language modeling in the continuous latent space. [46] use mel-spectrograms as the latent representation and model the per-step distribution with a regression loss. However, using a regression loss fails to capture the underlying distribution, requiring a post-AR refinement network and handcrafted repetition loss. [69] adopt a per-token diffusion loss [37], which requires iterative sampling for each autoregressive step, leading to significant inference latency. [73] models each step via flow matching [40], which also encounters similar problems. To this end, [26] proposes a patch-based continuous autoregressive framework that couples a language model for inter-patch prediction with a DiT [52] for intra-patch generation, thereby reducing the number of sampling steps. [87, 39] model the data at each autoregressive step with a Gaussian distribution or GMM and employ a VAE to obtain the target distribution parameters for each step. These approaches impose overly strong constraints on the latent space, limiting the model’s expressive capacity.
@@ 5.3 Evaluation Settings
We evaluate zero-shot speech synthesis performance using LibriSpeech test-clean set, ensuring that none of the test speakers are included in the training data. Following [72, 46], we use samples with durations between 4 and 10 seconds, resulting in a 2.2-hour subset comprising 1,234 samples and 40 unique speakers. Since the LibriSpeech consists of case-insensitive text without punctuation, while the expected input for models trained on LibriHeavy is unnormalized, we use the test-clean set of LibriSpeech-PC [45] to evaluate SLED. The LibriSpeech-PC is nearly identical to LibriSpeech, but its text has been restored to include capitalization and punctuation. [^3] The test-clean set of Librispeech-PC excludes a small number of low-quality samples (203 out of 2,620) from the original Librispeech test-clean set, thus the test subset after applying the same protocol contains only 1,154 samples. The evaluation is conducted under two settings: 1) 3s Prefix as Prompt: The model is provided with the text transcription and the first 3 seconds of the utterance as the prompt, and it is expected to perform speech continuation. 2) Reference Utterance as Prompt: The model is given a reference utterance from the same speaker along with its transcription as the prompt. Using the text of the target utterance, the model is expected to synthesize speech that retains the characteristics of the reference speaker. For this experiment, we follow the configuration outlined in [5]. We first filter the samples in LibriSpeech test-clean set based on length and order them by sample ID. For each speaker, the $i$-th speech sample is synthesized using the (i1)(i-1)-th sample as the prompt, while the first sample is synthesized using the last sample from the same speaker as the prompt. For Offline and Streaming Inference experiments, we follow previous data settings and no prompt speech is provided. We primarily use the following metrics to assess the intelligibility and in-context learning capability.
@@ 5.4 Main Results
Zero-shot TTS Table 1 presents the results of zero-shot TTS. In both the speech continuation and reference utterance prompting tasks, SLED achieves a word accuracy surpassing the ground truth (1.59/1.51 vs. 1.78 in WER-C), demonstrating its modeling ability. Interestingly, we found that using a reference utterance as a prompt results in more accurate synthesis compared to using the prefix, showing the ability of in-context learning. Comparisons of SLED performance across different data scales are provided in App. A. We found that 1000 hours of speech is sufficient to achieve most of the generation and in-context learning capabilities. Further scaling the training data to 50,000 hours enhances these abilities even more. Notably, a comparison between SLED and VALL-E highlights key differences in autoregressive modeling approaches. Both models use Encodec as the latent encoder, but SLED performs modeling in the continuous domain, whereas VALL-E operates in the discrete domain. VALL-E-like hierarchical models require an extra non-autoregressive local model ($\approx$159M) to predict residual tokens. In contrast, SLED relies on a lightweight generative module ($\approx$35M) for continuous vector generation, achieving better parameter efficiency. Despite this, SLED outperforms VALL-E on all metrics. These results show the superiority of our continuous speech language modeling approach. Compared with other continuous approaches, SLED achieves better or similar quality while enjoys the advantage in generation efficiency. MELLE incorporates an NAR refinement, whereas SLED is purely autoregressive and can generate in a streaming manner. FELLE must perform multi‐step integration iterations along a predefined ODE path at each decoding step, while SLED only requires a single forward pass through its generative module. Moreover, at roughly the same parameter scale, current speech language modeling approaches still exhibit a significant gap in voice cloning compared with traditional TTS models. However, Llasa [81] shows that scaling up can dramatically strengthen discrete autoregressive speech models, which inspires us to further explore the scaling capabilities of continuous autoregressive modeling approaches.
@@ 6.2 Effects of Classifier-free Guidance
6.3 Efficiency Analysis
We view SLED and DiTAR [26] as complementary approaches to more efficient continuous autoregressive speech modeling. DiTAR reduces the number of sequential steps via a semi-autoregressive framework, yielding time complexity between fully autoregressive and fully non-autoregressive, whereas SLED accelerates sampling in the continuous latent space at each autoregressive step. We compare the computational efficiency of the two methods in Table 4. All the metrics are evaluated by inferring 10 seconds of audio.
Table 4: Comparison of decoding efficiency between SLED and DiTAR [26].
Model #Params #Transformer Layer RTF (bs=1) FLOPs
SLED 0.2B 12 0.8 280G
DiTAR (patch size 4) 0.6B 6+36+6 0.66 2750G
Although DiTAR’s patching scheme reduces the number of sequential steps, the approach is hierarchical. Each step in DiTAR bears a heavy generation workload and therefore relies on a relatively large local decoder (DiT), in contrast to SLED’s lightweight MLPs for per-step generation. This burden can erode the nominal benefits of patching and, by iterating within a local DiT, may also compromise global consistency; DiTAR paper reports that the local-DiT module must incorporate historical-patch dependencies to generate correctly, increasing complexity and narrowing the method’s computational advantage. Empirically, SLED achieves a comparable RTF to DiTAR while using nearly 10× fewer FLOPs and roughly one-third the parameters.
@@ 7 Conclusion and Limitation
Acknowledgement
We gratefully acknowledge all the reviewers for their valuable comments and suggestions. This work was supported by the Natural Science Foundation of Beijing, China (Grant No. L257006).
@@ Appendix C More Analysis of Inference Efficiency
Appendix D Broader Impact
The speech language modeling method presented in this paper enables voice cloning, and the experimental results demonstrate that voice cloning is feasible to some extent. However, as the findings indicate, there remains a noticeable gap between the cloned voice and the original speaker’s voice. As such, a speaker verification model can be used to detect potential voice forgeries.